# VILA2: VILA Augmented VILA

### v1

@@ 4 Related Work
Limitations. Due to source constraints, we concentrate on the design of a self-augmented data curation pipeline and verify the 7B, 8B, and 40B models with 50M data. Larger models and more data can have the potential to lead to better VLM capabilities with self-augmenting abilities. We will address these aspects in future work.
@@ 5 Conclusions
This work has explored the techniques, insights, and benefits of using VLMs to self-improve its pretraining. We introduced two primary augmentation loops, one leveraging VLM’s general captioning capacities and the other harnessing their strength in specialized visual tasks. We demonstrated the feasibility of three ‘free lunch’ rounds for VLMs through self-bootstrapping, with further improvements via knowledge distillation from specialist VLMs. Our new VILA2 models demonstrate SOTA performances across a comprehensive set of benchmarks. Fruitful future directions include a deeper delve into the potential synergy between synthetic and real data toward training stronger foundation models.
@@ A.2 Specialist Acquisition
Specialty and expertise can be obtained via gathering existing data from the open-source community, human labeling, and annotating by domain-specific models, e.g. OCR models for text recognition and detection models for bounding box prediction. We also experimented with using open-world detectors like OWLv2[76] to automatically label bounding boxes, VLMs to generate detailed captions, and LLMs such as Llama3-70B-Instruct to merge the information for the grounded narration specialist. However, we found that language models introduced more hallucinations into the merged grounded narration. This is because many different detection labels can share the same meaning and refer to the same instance, making it difficult for the language model to perform the bipartite matching between bounding boxes and their text correspondence.
@@ A.3 SFT Data
Captioning MSR-VTT [77], TextCaps, Image Paragraph Captioning [78], LLaVAR, ShareGPT4V-100K
Reasoning CLEVR [79], NLVR, VisualMRC [80]
Multi-images ActivityNet-QA [81], iVQA [82], MSRVTT-QA, MSVD-QA [83], STEM-QA [84]
OCR DVQA, OCRVQA, ST-VQA [85], SynthDoG-en, TextOCR-GPT4V, ArxivQA
World Knowledge WIT [86]
General VQA ViQuAE [87], VQAv2-train, Visual Dialog [88], GQA-train [51], ScienceQA-train, SHERLOCK [89], Geo170K [90], MMC-Instruction [91], LRV-Instruction [92], RefCOCO-train [93]
Text-only FLAN-1M [94], MathInstruct [95], Dolly [96], GSM8K-ScRel-SFT [97]
@@
A.4 Training Detail
@@
  • Stage 2: Pre-training Stage. We gather a total of 51 million images, consisting of 25 million image-text pairs with the highest CLIP scores from COYO-700M, 25 million images in an interleaved image-text format from the MMC4-Core subset, and 1 million images with detailed captions from ShareGPT4V-Pretrain. During this stage, we unfreeze both the multi-modal projector and the language model to enhance comprehension of the diverse multi-modal training data. We use the augmented data here to replace the original COYO captions.
  • Stage 3: Supervised Fine-tuning Stage. After stage 2, we collect diverse visual question-answer pairs and unfreeze all parameters to fine-tune the model for general-purpose VQA capacities.
A.4.2 40B Model
For the VILA2-40B model, we skip the cost-intensive stage 2 and train the model with 7.5 million images randomly sampled from the 25 million COYO subset pairing with various caption sources: 2.5 million with original COYO captions, 2.5 million with $\text{VILA}_{3}$ re-captioned descriptions, and 2.5 million with $\text{VILA}_{3}$ spatial specialist re-captioned descriptions. Both the multi-modal projector and the language model remain unfrozen. Note that adding interleaved data, such as MMC4, can further boost the performance and we leave this potential to future work.
A.5 Hyperparameters
@@
A.6 Additional Re-caption Comparisons
We provide additional VILA2 recaptioned examples from the SAM dataset comparing the baseline captions from alternative methods, e.g., the widely adopted InstructBLIP Flan-T5 XXL [59] and LLaVA-NeXT-34B [36], with captions generated through various rounds of VILA2’s intermediate models in Figure 6 through Figure 8. The correct facts are marked with green, hallucinations are marked with red and spatial related information are marked with blue. From these examples, we can see that while the baseline caption is descriptive, it contains more hallucinations. As the re-captioning and training cycle progresses, the model tends to generate captions with more details and fewer hallucinations. Additionally, the spatial specialist introduces more spatial information into the recaptions, as expected.
Figure 6: More examples of comparison among captions from generation rounds of VILA2 and the widely used caption baseline of InstructBLIP [59].
Figure 7: More examples of comparison among captions from generation rounds of VILA2 and the widely used caption baseline of InstructBLIP [59].
Figure 8: More examples of comparison among captions from generation rounds of VILA2 and an additional caption baseline of LLaVA-NeXT [36]. LLaVA-NeXT-34B tends to generate longer caption with more hallucinations.

### v2

@@ 4 Related Work & Limitations
Limitations. Due to resource constraints, we concentrate on the design of a self-augmented data curation pipeline and verify the 7B, 8B, and 40B models with 51M data. Larger models (e.g., 70B and 405B) and more data (>0.5B pretrain data) can have the potential to lead to better VLM capabilities with self-augmenting abilities. We will address these aspects in future work.
@@ 5 Conclusions
This work has explored the techniques, insights, and benefits of using VLMs to self-improve its pretraining. We introduced two primary augmentation loops, one leveraging VLM’s general captioning capacities and the other harnessing their strength in specialized visual tasks. We demonstrated the feasibility of three ‘free lunch’ rounds for VLMs through self-bootstrapping, with further improvements via knowledge distillation from specialist VLMs. Our new VILA2 models demonstrate SOTA performances across a comprehensive set of benchmarks. Fruitful future directions include a deeper delve into the potential synergy between synthetic and real data to train stronger foundation models.
@@ A.2 Specialist Acquisition
Specialty and expertise can be obtained via gathering existing data from the open-source community, human labeling, and annotating by domain-specific models, e.g. OCR models for text recognition and detection models for bounding box prediction. We also experimented with using open-world detectors like OWLv2(Minderer et al., 2024) to automatically label bounding boxes, VLMs to generate detailed captions, and LLMs such as Llama3-70B-Instruct to merge the information for the grounded narration specialist. However, we found that language models introduced more hallucinations into the merged grounded narration. This is because many different detection labels can share the same meaning and refer to the same instance, making it difficult for the language model to perform the bipartite matching between bounding boxes and their text correspondence.
@@ A.3 SFT Data
Captioning MSR-VTT (Xu et al., 2016), TextCaps, LLaVAR, Image Paragraph Captioning (Krause et al., 2017), ShareGPT4V-100K
Reasoning CLEVR (Johnson et al., 2017), NLVR, VisualMRC (Tanaka et al., 2021)
Multi-images ActivityNet-QA (Yu et al., 2019), VQAv2-train, iVQA (Yang et al., 2021), MSRVTT-QA, STEM-QA (Shen et al., 2024)
OCR DVQA, OCRVQA, ST-VQA (Biten et al., 2019), SynthDoG-en, TextOCR-GPT4V, ArxivQA
World Knowledge WIT (Srinivasan et al., 2021)
General VQA ScienceQA-train, VQAv2-train, ViQuAE (Lerner et al., 2022), Visual Dialog (Das et al., 2017), GQA-train (Hudson & Manning, 2019), SHERLOCK (Hessel et al., 2022), Geo170K (Gao et al., 2023), MMC-Instruction (Liu et al., 2024b), LRV-Instruction (Liu et al., 2024a), RefCOCO-train (Yu et al., 2016)
Text-only FLAN-1M (Wei et al., 2022), MathInstruct (Yue et al., 2023), Dolly (Conover et al., 2023), GSM8K-ScRel-SFT (Yuan et al., 2023)
@@
A.4 Specialist Data
  1. 1.
    Spatial Specialists. We continued training the specialist from the stage 2, ALLaVA caption (Chen et al., 2024), and GPT-4V caption from ShareGPT4V (Chen et al., 2023b).
  2. 2.
    Grounding Specialist. We split the cleaned 4M grounded narration into Grounding-Short 3M and Grounding-Long 838K for a two-stage training process. In stage 3, we combined Grounding-Short 3M with ALLaVA caption (Chen et al., 2024) to adapt to new tasks of grounded narration while maintaining the narrative ability. In stage 4, we combine Grounding-Long 838K with Shikra GPT-4 (Chen et al., 2023a), Visual7W (Zhu et al., 2016), LLaVA-SFT, and 100K GPT-4V captions from ShareGPT4V to sustain both narrative and instruction following capacities.
  3. 3.
    OCR Specialist. We trained our OCR specialist with various internet datasets focused on text recognition, understanding, and reasoning, including LLaVA-SFT, TextOCR-GPT4V (Carter, 2024), SynthDoG-En (Kim et al., 2022), OCRVQA (Anand et al., 2019), TextCaps (Sidorov et al., 2020), ArxivQA (Li et al., 2024a), DocVQA (Kafle et al., 2018), AI2D (Kembhavi et al., 2016), ChartQA (Masry et al., 2022), LLaVAR (Zhang et al., 2023) and 35 OCR-related datasets from The Cauldron (Laurençon et al., 2024).
A.5 Training Detail
@@
  • Stage 2: Pretraining Stage. We gather a total of 51 million images, consisting of 25 million image-text pairs with the highest CLIP scores from COYO-700M, 25 million images in an interleaved image-text format from the MMC4-Core subset, and 1 million images with detailed captions from ShareGPT4V-Pretrain. During this stage, we unfreeze both the multi-modal projector and the language model to enhance comprehension of the diverse multi-modal training data. We use the augmented data here to replace the original COYO captions.
  • Stage 3: Supervised Finetuning Stage. After stage 2, we collect diverse visual question-answer pairs and unfreeze all parameters to finetune the model for general-purpose VQA capacities.
A.5.2 40B Model
For the VILA2-40B model, we skip the cost-intensive stage 2 and train the model with 7.5 million images randomly sampled from the 25 million COYO subset pairing with various caption sources: 2.5 million with original COYO captions, 2.5 million with $\text{VILA}_{3}$ re-captioned descriptions, and 2.5 million with $\text{VILA}_{3}$ spatial specialist re-captioned descriptions. Both the multi-modal projector and the language model remain unfrozen. Note that adding interleaved data, such as MMC4, can further boost the performance and we leave this potential to future work. A detailed profiling of 40B performances over benchmarks is also included as Table 12.
[…]
VILA2-40B (ours) Yi-34B 448 85.1 64.7 62.2 93.2 75.9 83.9 82.9 77.0 93.6 53.4
Table 12: Improvements from 8B to 40B checkpoints on 10 visual-language benchmarks.
A.6 Hyperparameters
@@
A.7 Additional Re-caption Comparisons
We provide additional VILA2 recaptioned examples from the SAM dataset comparing the baseline captions from alternative methods, e.g., the widely adopted InstructBLIP Flan-T5 XXL (Dai et al., 2023) and LLaVA-NeXT-34B (Liu et al., 2024c), with captions generated through various rounds of VILA2’s intermediate models in Figure 6 through Figure 8. The correct facts are marked with green, hallucinations are marked with red and spatial related information are marked with blue. From these examples, we can see that while the baseline caption is descriptive, it contains more hallucinations. As the re-captioning and training cycle progresses, the model tends to generate captions with more details and fewer hallucinations. Additionally, the spatial specialist introduces more spatial information into the recaptions, as expected.
Figure 6: More examples of comparison among captions from generation rounds of VILA2 and the widely used caption baseline of InstructBLIP (Dai et al., 2023).
Figure 7: More examples of comparison among captions from generation rounds of VILA2 and the widely used caption baseline of InstructBLIP (Dai et al., 2023).
Figure 8: More examples of comparison among captions from generation rounds of VILA2 and an additional caption baseline of LLaVA-NeXT (Liu et al., 2024c). LLaVA-NeXT-34B tends to generate longer caption with more hallucinations.